Can someone show me how to evaluate the values of x giving as 1,2,4,8 and 16 for the equation f(x)=logx Can someone show me how to work this problem for one of the values so that I can continue with the remaining values for x. Thank you.
Answer by Alan3354(69443)

Can someone show me how to evaluate the values of x giving as 1,2,4,8 and 16 for the equation f(x)=logx
f(1) = log(1) = 0
f(2) = log(2) =~0.30103
f(4) = log(4) = 2*log(2) 4 = 2^2
f(8) = log(8) = 3*log(2) 8 = 2^3
f(16) = 4*log(2) 16 = 2^4
etc.